# AppReferences

### v1

@@ 3.1 Causal Estimands
In this paper, we introduce a formal and novel definition of principal strata that takes into account the continuous nature of the post-treatment variable.
We assume that the units can be marginally—respect to the treatment level—grouped into clusters with finite first moments. Consequently, each unit $i$ belongs to either: the dissociative stratum if $S_{i}(1)=S_{i}(0)$, which is equivalent to $\mathbb{E}_{S_{i}^{(1)}}[P_{i}(1)]=\mathbb{E}_{S_{i}^{(0)}}[P_{i}(0)]$; or the associative positive stratum if $S_{i}(1)\not=S_{i}(0)$ and $\mathbb{E}_{S_{i}^{(1)}}[P_{i}(1)]>\mathbb{E}_{S_{i}^{(0)}}[P_{i}(0)]$ (now on, we indicate these conditions as $S_{i}(1)\succ S_{i}(0)$); or the associative negative stratum if $S_{i}(1)\not=S_{i}(0)$ and $\mathbb{E}_{S_{i}^{(1)}}[P_{i}(1)]<\mathbb{E}_{S_{i}^{(0)}}[P_{i}(0)]$ (now on, we indicate these conditions as $S_{i}(1)\prec S_{i}(0)$).
Definition 1 (Causal Estimands)
Conditionally on strata allocation, we can define the causal estimands for the expected dissociative effect (EDE) and the expected associative effects (EAE) as
$EDE=\tau_{0}=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid S_{i}(1)=S_{i}(0)],$
$EAE_{+}=\tau_{+1}=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid S_{i}(1)\succ S_{i}(0)],$
$EAE_{-}=\tau_{-1}=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid S_{i}(1)\prec S_{i}(0)].$
These causal estimands are not directly identifiable from the data due to the conditioning on unobservable quantities (i.e., $S_{i}(1),S_{i}(0)$). However, they can be weak identified invoking the causal assumptions introduced above—i.e., SUTVA and strongly ignorable treatment assignment.
Theorem 1
Invoking assumptions (1) and (2), the expected value of each potential outcome can be rewritten as the following
$\mathbb{E}\big{[}Y_{i}(t)\mid S_{i}(1)=s_{1},S_{i}(0)=s_{0},X_{i}=x\big{]}$
@@
×$\quad\quad\times Pr\big{(}P_{i}(1)=p_{1},P_{i}(0)=p_{0}\mid T_{i}=t,S_{i}(1)=s_{1},S_{i}(0)=s_{0},X_{i}=x\big{)}dp_{0}p_{1}dx;$ (8)
where the inner expectations $\mathbb{E}\big{[}Y_{i}\mid T_{i}=t,X_{i}=x,P_{i}(1)=p_{1},P_{i}(0)=p_{0}\big{]}$ is estimate with the outcome model $Y_{i}\mid T_{i},X_{i},P_{i}(1),P_{i}(0)$, as well the probability $\Pr\big{(}P_{i}(1)=p_{1},P_{i}(0)=p_{0}\mid T_{i}=t,S_{i}(1)=s_{1},S_{i}(0)=s_{0},X_{i}=x\big{)}$, that is defined by the model for the potential post-treatment variables.
The proof is reported in Appendix B.
@@ 3.2 Discovery of Principal Strata
The definition of the previous (3) for the post-treatment variable model allows us to characterize two distinct specific mixtures for each of the two treatment levels $t=\{0,1\}$, taking into account the different heterogeneity in the potential post-treatment variables induced by the covariates $X$. Simultaneously, the parameters of the shared random kernel $\{\psi_{l}\}_{l\geq 1}$ induce desirable characteristics in the potential post-treatment variables, such as (i) the potential post-treatment variables $\{P(1),P(0)\}$ are not independent, even conditionally, of the covariates $X$—this is clear from the posterior distribution reported in Section 3.3 and Appendix D—and (ii) there is a nonnull probability that any $i$ belongs to the defined dissociative stratum—see Theorem 2.
The fundamental element in explaining this is the definition of the latent categorical variables $\{S_{i}^{(0)},S_{i}^{(1)}\}$ for each unit $i\in\{1,\dots,n\}$, which describe the probability that each unit will be allocated in the components of the mixture. The Bayesian nonparametric mixtures are well-known for their ability to capture and describe the cluster allocation of the observations; however, we focus on the causal inference goal to describe the principal strata.
In our fully Bayesian approach, pairs $\{S_{i}^{(0)},S_{i}^{(1)}\}$ are random variables that define the two marginal distributions of the random unit partition (Quintana, 2006), based on the two potential post-treatment variables. However, we want to define the join distribution of them, specifically the unit allocation in the principal strata. This join distribution is not directly observed from the data, due to missing data problems, but we can estimate this thanks to the characteristic structure of the proposed CASBAH model.
The proposed CASBAH model is defined such that the atoms $\{\psi_{l}\}_{l\geq 1}$ are shared between the two potential outcomes of the post-treatment variable of the same unit. Therefore, the model informs us about the posterior probability that $\{S_{i}^{(0)},S_{i}^{(1)}\}$ is associated with the same or different atoms, for each unit $i$. For example, the probability for the unit $i\in\{1,\dots,n\}$ to belong to the dissociative stratum—i.e., $S_{i}(0)=S_{i}(1)$—is equivalent to the probability that the atom for the control is the same for the treated. Such a probability is not null and can be derived, a priori, by means of the following theorem.
Theorem 2
Given a unit $i\in\{1,\dots,n\}$, with covariates $x_{i}$, following the (2) distribution, the probability to belong to the dissociative stratum for each unit $i$ is a function of its covariates $x_{i}$, defined as following
$\Pr[S_{i}(0)=S_{i}(1)]=\frac{\rho_{2}(x_{i})[(1+\rho_{2}(x_{i})-2\rho_{1}(x_{i}))^{L}-1]}{\rho_{2}(x_{i})-2\rho_{1}(x_{i})}$ (9)
where $\rho_{1}(x_{i})=\mathbb{E}\big{[}\Phi\big{(}\alpha(x_{i})\big{)}\big{]}$ and $\rho_{2}(x_{i})=\mathbb{E}\big{[}\Phi\big{(}\alpha(x_{i})\big{)}^{2}\big{]}$ and $0\leq\Pr[S_{i}(0)=S_{i}(1)]\leq 1$.
The proof is reported in Appendix C.
@@ 3.3 Posterior Inference
As Theorem 1 shows, the principal causal estimand are functions of the models for potential outcomes and the potential post-treatment variable that involve parameters in both models. Following the Bayesian paradigm, the estimation is based on the posterior distribution of the parameters conditional on the data.
Sampling from the posterior joint distribution is straightforward via Gibbs sampling. In particular, the algorithm—described in detail in Section D—takes inspiration from those proposed by Teh et al. (2004, 2006); Teh and Jordan (2009), for the hierarchical structure, and by Fasano et al. (2022) for the probit regression in the weights.
Assuming a general linear regression for the outcome model that depends on the parameter $\theta_{0}^{Y}$ for the outcome under control and $\theta_{1}^{Y}$ for the treated outcome, the Gibbs sampler interests the following distributions:
@@
    $\Pr\{S_{i}^{(t)}=l\}\propto\pi_{l}^{(t)}({x}_{i})\Pr(P_{i}(t)\mid S_{i}^{(t)}=l,X_{i};\eta_{l},\sigma_{l}),$
    where $\pi_{l}^{(t)}({x}_{i})=\Phi(\alpha_{l}^{(t)}({x}_{i}))\prod_{r<l}(1-\Phi(\alpha_{r}^{(t)}({x}_{i})))p$ for $l=1,\dots,L-1$ and $\Phi(\alpha_{L}^{(t)}({x}_{i}))=1$.
@@
    following (Fasano et al., 2022) demonstration—detail in Appendix A.
@@ 4 Simulation Study
The performance of the proposed CASBAH mixture model is evaluated by means of a simulation study. Our objective is to investigate the model’s ability to (i) accurately impute the missing post-treatment and outcome variables, that is, to assess the potential bias of the expected value $E[P_{i}(1)-P_{i}(0)]$ and $E[Y_{i}(1)-Y_{i}(0)]$ over the units $i=1\dots,n$—, (ii) correctly identify the principal strata. To achieve this, we conducted simulations under five different data-generating mechanisms and analyzed the results to understand the model’s behavior in different scenarios.
Specifically, we assume a linear regression model for the outcome model, defined as follows:
@@
with $\mu_{\theta}=0$, $\sigma_{\theta}=10$, $\mu_{\lambda}=0$, and $\sigma_{\lambda}=2$. Clearly, our model can accommodate more complex choices for the outcome, but we decided to opt for a simpler model following what had already been done in the literature by Schwartz et al. (2011). The performance of the proposed approach is compared to those obtained with the Schwartz et al. (2011)’s model (hereafter referred to as SLM).
@@
Each scenario assumes a different conformation of the strata for the continuous post-treatment variable $P_{i}=(P_{i}(0),P_{i}(1))\in\mathbb{R}^{2}$. Each stratum is obtained by introducing, for each unit, categorical variables $S_{i}(0)$ and $S_{i}(1)$, for control and treatment levels, respectively, with the vector of probabilities that depend on the values of the confounders and allocates the unit in different clusters. Conditionally on the cluster allocation $S_{i}^{(t)}=s$, with $s$ that has the same support for control and treatment levels, we simulate both potential post-treatment variables—under control and under treatment, respectively—as
$P_{i}(0)|S_{i}(0)=s\sim\mathcal{N}(\eta_{s},\sigma^{2}_{s}),\quad P_{i}(1)|S_{i}(1)=s\sim\mathcal{N}(\eta_{s},\sigma^{2}_{s}).$
@@
Scenario 1: We investigate a situation in which there are two strata: one with a dissociative effect and one with a positive associative effect. In particular, for $S_{i}(0)=S_{i}(1)=1$ $P_{i}(0),P_{i}(1)\sim\mathcal{N}(1,0.05)$, and for $S_{i}(0)=2$ and $S_{i}(1)=3$ $P_{i}(0)\sim\mathcal{N}(2,0.05)$ and $P_{i}(1)\sim\mathcal{N}(3,0.05)$. The regression parameters for the Y-model are $\theta^{(0)}=(1,2)$ and $\theta^{(1)}=(1,2,-1,0.5)$.
Scenario 2: We focus on a case where we have a dissociative stratum—for $S_{i}(0)=S_{i}(1)=1$ both $P(0)$ and $P(1)$ are simulated from $\mathcal{N}(2,0.05)$—, an associative stratum with a positive effect—for $S_{i}(0)=2$ and $S_{i}(1)=3$ $P_{i}(0)\sim\mathcal{N}(2,0.05)$ and $P_{i}(1)\sim\mathcal{N}(3,0.05)$—, and dissociative stratum with a negative effect—for $S_{i}(0)=2$ and $S_{i}(1)=1$ $P_{i}(0)\sim\mathcal{N}(2,0.05)$ and $P_{i}(1)\sim\mathcal{N}(1,0.05)$. The regression parameters for Y-model are $\theta^{(0)}=(1,2)$ and $\theta^{(1)}=(1,1.2,-1,1)$.
Scenario 3: This scenario corresponds to Scenario 1 with closer atoms for the strata and different variances. In particular, the dissociative stratum has $(P_{i}(0)|S_{i}(0)=1)=(P_{i}(1)|S_{i}(1)=1)\sim\mathcal{N}(1.5,0.12)$, and the associative stratum has $(P_{i}(0)|S_{i}(0)=2)\sim\mathcal{N}(2,0.1)$ and $(P_{i}(1)|S_{i}(1)=3)\sim\mathcal{N}(2.5,0.08)$. The regression parameters for Y-model are $\theta^{(0)}=(1,2)$ and $\theta^{(1)}=(1,1.2,-0.8,0.5)$.
Scenario 4: This scenario corresponds to Scenario 2 with closer atoms for the strata and different variances. In particular, the dissociative stratum has $(P_{i}(0)|S_{i}(0)=1)=(P_{i}(1)|S_{i}(1)=1)\sim\mathcal{N}(1.5,0.12)$, the associative positive stratum has $(P_{i}(0)|S_{i}(0)=2)\sim\mathcal{N}(2,0.1)$ and $(P_{i}(1)|S_{i}(1)=3)\sim\mathcal{N}(2.5,0.08)$, and the associative negative stratum has $(P_{i}(0)|S_{i}(0)=2)\sim\mathcal{N}(2,0.1)$ and $(P_{i}(1)|S_{i}(1)=1)\sim\mathcal{N}(1.5,0.12)$. The regression parameters for Y-model are $\theta^{(0)}=(1,2)$ and $\theta^{(1)}=(1,1.2,-0.8,0.5)$.
Scenario 5: We investigate the scenario with the three strata when the number of confounders increases, in particular the treatment variable and cluster allocation variables that depend on five confounders. The dissociative stratum has $(P_{i}(0)|S_{i}(0)=1)=(P_{i}(1)|S_{i}(1)=1)\sim\mathcal{N}(2,0.05)$, the associative positive stratum has $(P_{i}(0)|S_{i}(0)=3)\sim\mathcal{N}(3,0.05)$ and $(P_{i}(1)|S_{i}(1)=4)\sim\mathcal{N}(4,0.05)$, and the associative negative stratum has $(P_{i}(0)|S_{i}(0)=2)\sim\mathcal{N}(2,0.05)$ and $(P_{i}(1)|S_{i}(1)=1)\sim\mathcal{N}(1,0.05)$. The regression parameters for Y-model are $\theta^{(0)}=(1,2)$ and $\theta^{(1)}=(1,1.2,-1,0.5)$.
@@
Table 1 reports the median and interquartile range (IQR) of the bias for the expected value of the posterior distribution of sample average of $P_{i}(1)-P_{i}(0)$ and $Y_{i}(1)-Y_{i}(0)$, for $i\in\{1,\dots,n\}$. The results for the five scenarios for the proposed model CASBAH show a good ability of the proposed model to impute the missing variables and capture the true distribution, indeed the medians of the bias are close to zero and the interquartile range is reasonable according to the simulated variability. The comparison with the Schwartz et al. (2011)’s model, SLM, underlines the superiority of CASBAH, which obtains medians of the bias closer to zero and a more contained IQR for $\mathbb{E}[P_{i}(1)-P_{i}(0)]$ in all scenarios and almost of them for $\mathbb{E}[Y_{i}(1)-Y_{i}(0)]$.
Scenario 1 Scenario 2 Scenario 3 Scenario 4 Scenario 5
Post-treatment variable
CASBAH Median -0.0164 0.0027 -0.0318 0.0027 0.0011
IQR 0.0315 0.0254 0.0462 0.0325 0.0183
SLM Median -0.0637 0.3388 0.1254 0.3280 0.3572
IQR 0.0989 0.0986 0.0956 0.0963 0.1216
Outcome
CASBAH Median 0.0024 0.0256 0.0067 0.0419 0.0344
IQR 0.2235 0.1550 0.4661 0.1357 0.1316
SLM Median -1.9891 -1.8265 -1.4363 -1.3115 -1.8856
IQR 0.2546 0.2110 0.1904 0.1836 0.3727
@@
Table 1: Median and interquartile range (IQR) of the bias for the expected value of the posterior distribution of sample average of $P_{i}(1)-P_{i}(0)$ and $Y_{i}(1)-Y_{i}(0)$, for $i\in\{1,\dots,n\}$. Values reported for CASBAH and SLM.
@@
To evaluate the correct identification of the principal strata, we use the adjusted Rand index (ARI, introduced by Hubert and Arabie (1985)). The ARI values for the five scenarios are reported in Table 2. For all scenarios, the index is close to $1$ and confirms that the proposed CASBAH model can correctly identify the principal strata and, combined with the good missing data imputation, allows us to estimate the expected associative and dissociative effects. The SLM model does not identify the principal strata according to our definition or simultaneously to the model estimation, in opposition to CASBAH.
Scenario 1 Scenario 2 Scenario 3 Scenario 4 Scenario 5
Mean 0.9850 0.9906 0.9706 0.9717 0.9154
SD 0.0997 0.0597 0.1021 0.0892 0.1577
@@
Table 2: Adjusted rand index for the five simulated scenarios computed on the point estimated partitions obtained with the proposed model CASBAH. mean and empirical standard deviation (sd) are reported.

### v2

@@ 3.1 Causal Estimands
To define our causal estimands of interest, we postulate that the units can be marginally grouped into clusters with finite first moments with respect to the treatment level. Consequently, each unit $i$ can belong to: (i) the dissociative stratum if $S_{i}^{(1)}=S_{i}^{(0)}$, which is equivalent to $\mathbb{E}_{S_{i}^{(1)}}[P_{i}(1)]=\mathbb{E}_{S_{i}^{(0)}}[P_{i}(0)]$; or (ii) the associative positive stratum if $S_{i}^{(1)}\not=S_{i}^{(0)}$ and $\mathbb{E}_{S_{i}^{(1)}}[P_{i}(1)]>\mathbb{E}_{S_{i}^{(0)}}[P_{i}(0)]$ (throughout the paper, we indicate this condition as $S_{i}^{(1)}\succ S_{i}^{(0)}$); or (iii) the associative negative stratum if $S_{i}^{(1)}\not=S_{i}^{(0)}$ and $\mathbb{E}_{S_{i}^{(1)}}[P_{i}(1)]<\mathbb{E}_{S_{i}^{(0)}}[P_{i}(0)]$ (throughout, we indicate this condition as $S_{i}^{(1)}\prec S_{i}^{(0)}$).
Definition 3.1 (Causal Estimands).
Conditionally on the strata allocation, we define the causal estimands for the expected dissociative effect (EDE) and the expected associative effects (EAE) as
$EDE=\tau_{0}=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid S_{i}^{(1)}=S_{i}^{(0)}],$
$EAE_{+}=\tau_{+1}=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid S_{i}^{(1)}\succ S_{i}^{(0)}],$
$EAE_{-}=\tau_{-1}=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid S_{i}^{(1)}\prec S_{i}^{(0)}].$
These causal estimands are not directly identifiable from the data due to the conditioning on unobservable quantities (i.e., $S_{i}^{(1)},S_{i}^{(0)}$). However, they can be weakly identified by invoking Assumptions (1) and (2).
Theorem 3.2.
Invoking assumptions (2.1) and (2.1), the expected value of each potential outcome can be rewritten as the following
$\mathbb{E}\big{[}Y_{i}(t)\mid S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big{]}$
@@
×$\quad\quad\times Pr\big{(}P_{i}(1)=p_{1},P_{i}(0)=p_{0}\mid T_{i}=t,S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big{)}dp_{0}p_{1}dx;$ (8)
where the inner expectations $\mathbb{E}\big{[}Y_{i}\mid T_{i}=t,X_{i}=x,P_{i}(1)=p_{1},P_{i}(0)=p_{0}\big{]}$ is estimate with the outcome model $Y_{i}\mid T_{i},X_{i},P_{i}(1),P_{i}(0)$, as well the probability $\Pr\big{(}P_{i}(1)=p_{1},P_{i}(0)=p_{0}\mid T_{i}=t,S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big{)}$, that is defined by the model for the potential post-treatment variables.
The proof is reported in the Supplementary Material.
@@ 3.2 Discovery of Principal Strata
While Bayesian nonparametric mixtures excel at capturing cluster structures, our focus shifts to leveraging this framework for causal inference. To estimate our causal estimands of interest, we need to characterize the joint distribution of $\{S_{i}^{(0)},S_{i}^{(1)}\}$ so that we can determine each unit’s allocation to a principal strata. This joint distribution cannot be directly observed due to the fundamental problem of causal inference. Nevertheless, our proposed Bayesian nonparametric model allows for robust estimation of this joint distribution through its hierarchical construction and shared parameters, effectively addressing this problem. In this section, we illustrate how our proposed Bayesian nonparametric model enables the characterization of principal strata in a statistically robust and fully data-adaptive manner.
We start from noting that the latent categorical variables $\{S_{i}^{(0)},S_{i}^{(1)}\}$ for each unit $i\in\{1,\ldots,n\}$ serve a dual purpose. First, they determine the allocation probabilities to specific mixture components. Second, they simultaneously define the marginal distributions of the random unit partition based on potential post-treatment outcomes (Quintana, 2006). This formulation creates a natural bridge between mixture model clustering and principal stratification, allowing us to characterize treatment effect heterogeneity through a principled probabilistic structure.
Our proposed model is defined such that the atoms $\{\psi_{l}\}_{l\geq 1}$ are shared between the two potential post-treatment outcomes for the same unit. This formulation yields precise posterior probabilities that quantify, for each unit $i$, whether the latent indicators $\{S_{i}^{(0)},S_{i}^{(1)}\}$ map to identical or distinct mixture components. Specifically, for any unit $i\in\{1,\ldots,n\}$, the probability of membership in the dissociative stratum—where $S_{i}^{(0)}\neq S_{i}^{(1)}$—equals the probability that the atom under control differs from the atom under treatment. This probability is provably non-zero and can be derived a priori through the following theorem.
Theorem 3.3.
Given a unit $i\in\{1,\dots,n\}$, with covariates $x_{i}$ following (2), the probability to belong to the dissociative stratum for each unit $i$ is a function of its covariates $x_{i}$, defined as:
$\Pr[S_{i}^{(0)}=S_{i}^{(1)}]=\frac{\rho_{2}(x_{i})[(1+\rho_{2}(x_{i})-2\rho_{1}(x_{i}))^{L}-1]}{\rho_{2}(x_{i})-2\rho_{1}(x_{i})}$ (9)
where $\rho_{1}(x_{i})=\mathbb{E}\big{[}\Phi\big{(}\alpha(x_{i})\big{)}\big{]}$ and $\rho_{2}(x_{i})=\mathbb{E}\big{[}\Phi\big{(}\alpha(x_{i})\big{)}^{2}\big{]}$ and $0\leq\Pr[S_{i}^{(0)}=S_{i}^{(1)}]\leq 1$.
The proof is reported in the Supplementary Material. The representation of the probability to belong to the dissociative stratum in (9) is further illustrated in Figure 1.
[…]
@@ 3.3 Posterior Inference
Theorem 3.2 establishes that principal causal estimands are characterized as functions of both the primary response and potential post-treatment outcomes, each involving specific model parameters. Following the Bayesian paradigm, we estimate these parameters through their posterior distributions conditional on the observed data. This approach provides not only point estimates but also complete uncertainty quantification.
Sampling from the posterior joint distribution is straightforward via Gibbs sampling. In particular, the proposed algorithm (described in detail in the Supplementary Material) takes inspiration from Teh et al. (2004, 2006); Teh & Jordan (2009) for the hierarchical structure, and from Fasano et al. (2022) for the probit regression in the weights. We assume a general linear regression for the outcome model that depends on the parameter $\theta_{0}^{Y}$ for the outcome under control and $\theta_{1}^{Y}$ for the treated outcome.
The implementation of the Gibbs sampler proceeds through the following steps.
@@
    $\Pr\{S_{i}^{(t)}=l\}\propto\pi_{l}^{(t)}(\mathbf{x}_{i})\Pr(P_{i}(t)\mid S_{i}^{(t)}=l,X_{i};\eta_{l},\sigma_{l}),$
    where $\pi_{l}^{(t)}(\mathbf{x}_{i})=\Phi(\alpha_{l}^{(t)}(\mathbf{x}_{i}))\prod_{r<l}(1-\Phi(\alpha_{r}^{(t)}(\mathbf{x}_{i})))p$ for $l=1,\dots,L-1$ and $\Phi(\alpha_{L}^{(t)}(\mathbf{x}_{i}))=1$.
@@
    following (Fasano et al., 2022) demonstration, detail in the Supplementary Material.
@@ 4 Simulation Study
The performance of the proposed mixture model is evaluated through a simulation study. Our objective is to investigate the model’s ability to (i) accurately impute the missing post-treatment and outcome variables, that is, to assess the potential bias of the expected value $E[P_{i}(1)-P_{i}(0)]$ and $E[Y_{i}(1)-Y_{i}(0)]$ over the units $i=1\dots,n$, (ii) correctly identify the principal strata and estimate the principal causal effects. To achieve this, we conducted simulations under five different data-generating mechanisms and analyzed the results to understand the model’s behavior in different scenarios.
The performance of our proposed approach is compared to those obtained with the Schwartz et al. (2011)’s model and the copula model proposed by Lu et al. (2023).
Commonly across the scenarios, we assume a linear regression model for the outcome model, defined as follows:
@@
with $\mu_{\theta}=0$, $\sigma_{\theta}=10$, $\mu_{\lambda}=0$, and $\sigma_{\lambda}=2$. Clearly, our model can accommodate more complex choices for the outcome, but we decided to opt for a simpler model following what had already been done in the literature by Schwartz et al. (2011) and Lu et al. (2023).
@@
Each scenario assumes a different conformation of the strata for the continuous post-treatment variable $P_{i}=(P_{i}(0),P_{i}(1))\in\mathbb{R}^{2}$. Each stratum is obtained by introducing, for each unit, categorical variables $S_{i}^{(0)}$ and $S_{i}^{(1)}$, for control and treatment levels, respectively, with the vector of probabilities that depend on the values of the confounders and allocates the unit in different clusters. Conditionally on the cluster allocation $S_{i}^{(t)}=s$, with $s$ that has the same support for control and treatment levels, we simulate both potential post-treatment variables, under control and under treatment, respectively, as
$P_{i}(0)|S_{i}^{(0)}=s\sim\mathcal{N}(\eta_{s},\sigma^{2}_{s}),\quad P_{i}(1)|S_{i}^{(1)}=s\sim\mathcal{N}(\eta_{s},\sigma^{2}_{s}).$
@@
Scenario 1: We investigate a situation in which there are two strata: one with a dissociative effect and one with a positive associative effect. In particular, for $S_{i}^{(0)}=S_{i}^{(1)}=1$ $P_{i}(0),P_{i}(1)\sim\mathcal{N}(1,0.05)$, and for $S_{i}^{(0)}=2$ and $S_{i}^{(1)}=3$ $P_{i}(0)\sim\mathcal{N}(2,0.05)$ and $P_{i}(1)\sim\mathcal{N}(3,0.05)$. The regression parameters for the Y-model are $\theta^{(0)}=(1,2)$ and $\theta^{(1)}=(1,2,-1,0.5)$.
Scenario 2: We focus on a case where we have a dissociative stratum, for $S_{i}^{(0)}=S_{i}^{(1)}=1$ both $P(0)$ and $P(1)$ are simulated from $\mathcal{N}(2,0.05)$, an associative stratum with a positive effect, for $S_{i}^{(0)}=2$ and $S_{i}^{(1)}=3$ $P_{i}(0)\sim\mathcal{N}(2,0.05)$ and $P_{i}(1)\sim\mathcal{N}(3,0.05)$, and dissociative stratum with a negative effect, for $S_{i}^{(0)}=2$ and $S_{i}^{(1)}=1$ $P_{i}(0)\sim\mathcal{N}(2,0.05)$ and $P_{i}(1)\sim\mathcal{N}(1,0.05)$. The regression parameters for Y-model are $\theta^{(0)}=(1,2)$ and $\theta^{(1)}=(1,1.2,-1,1)$.
Scenario 3: This scenario corresponds to Scenario 1 with closer atoms for the strata and different variances. In particular, the dissociative stratum has $(P_{i}(0)|S_{i}^{(0)}=1)=(P_{i}(1)|S_{i}^{(1)}=1)\sim\mathcal{N}(1.5,0.12)$, and the associative stratum has $(P_{i}(0)|S_{i}^{(0)}=2)\sim\mathcal{N}(2,0.1)$ and $(P_{i}(1)|S_{i}^{(1)}=3)\sim\mathcal{N}(2.5,0.08)$. The regression parameters for Y-model are $\theta^{(0)}=(1,2)$ and $\theta^{(1)}=(1,1.2,-0.8,0.5)$.
Scenario 4: This scenario corresponds to Scenario 2 with closer atoms for the strata and different variances. In particular, the dissociative stratum has $(P_{i}(0)|S_{i}^{(0)}=1)=(P_{i}(1)|S_{i}^{(1)}=1)\sim\mathcal{N}(1.5,0.12)$, the associative positive stratum has $(P_{i}(0)|S_{i}^{(0)}=2)\sim\mathcal{N}(2,0.1)$ and $(P_{i}(1)|S_{i}^{(1)}=3)\sim\mathcal{N}(2.5,0.08)$, and the associative negative stratum has $(P_{i}(0)|S_{i}^{(0)}=2)\sim\mathcal{N}(2,0.1)$ and $(P_{i}(1)|S_{i}^{(1)}=1)\sim\mathcal{N}(1.5,0.12)$. The regression parameters for Y-model are $\theta^{(0)}=(1,2)$ and $\theta^{(1)}=(1,1.2,-0.8,0.5)$.
Scenario 5: We investigate the scenario with the three strata when the number of confounders increases, in particular the treatment variable and cluster allocation variables that depend on five confounders. The dissociative stratum has $(P_{i}(0)|S_{i}^{(0)}=1)=(P_{i}(1)|S_{i}^{(1)}=1)\sim\mathcal{N}(2,0.05)$, the associative positive stratum has $(P_{i}(0)|S_{i}^{(0)}=3)\sim\mathcal{N}(3,0.05)$ and $(P_{i}(1)|S_{i}^{(1)}=4)\sim\mathcal{N}(4,0.05)$, and the associative negative stratum has $(P_{i}(0)|S_{i}^{(0)}=2)\sim\mathcal{N}(2,0.05)$ and $(P_{i}(1)|S_{i}^{(1)}=1)\sim\mathcal{N}(1,0.05)$. The regression parameters for Y-model are $\theta^{(0)}=(1,2)$ and $\theta^{(1)}=(1,1.2,-1,0.5)$.
@@
Table 4 reports the median and interquartile range of the bias for the expected value of the posterior distribution of the sample average of $P_{i}(1)-P_{i}(0)$ and $Y_{i}(1)-Y_{i}(0)$, for $i\in\{1,\dots,n\}$.
The results for the five scenarios demonstrate the strong ability of our proposed model to impute missing variables and accurately capture the true distribution across different data generating processes. The medians of the bias are close to zero and the interquartile range is reasonable given the simulated variability. In contrast, the two competing models highlight their limitations in recovering the same information. The algorithm from Lu et al. (2023) does not estimate the distribution of the potential outcome for the post-treatment variable, while the model from Schwartz et al. (2011) exhibits significant bias in four out of five scenarios and a higher interquartile range than our model across all scenarios. Furthermore, the comparison of $\mathbb{E}[Y_{i}(1)-Y_{i}(0)]$ reveals substantial bias in all five scenarios for both competing models Lu et al. (2023) and Schwartz et al. (2011).
[…]
Bias comparison of the two methods based on different simulation scenarios. Scenario 1 Scenario 2 Scenario 3 Scenario 4 Scenario 5 $\mathbb{E}[P(1)-P(0)]$ Our proposed method Median -0.0164 0.0027 -0.0318 0.0027 0.0011 IQR 0.0315 0.0254 0.0462 0.0325 0.0183 Schwartz et al. (2011) Median -0.0637 0.3388 0.1254 0.3280 0.3572 IQR 0.0989 0.0986 0.0956 0.0963 0.1216 Lu et al. (2024) Median IQR $\mathbb{E}[Y(1)-Y(0)]$ Our proposed method Median 0.0024 0.0256 0.0067 0.0419 0.0344 IQR 0.2235 0.1550 0.4661 0.1357 0.1316 Schwartz et al. (2011) Median -1.9891 -1.8265 -1.4363 -1.3115 -1.8856 IQR 0.2546 0.2110 0.1904 0.1836 0.3727 Lu et al. (2024) Median -1.3491 0.3070 -0.5848 0.1393 0.0751 IQR 0.1092 0.1206 0.1009 0.0869 0.1078 {tabnote} IQR: interquartile range.
@@
To assess the accurate identification of the principal strata, we use the adjusted Rand index (Hubert & Arabie, 1985) and evaluate the estimation of the principal causal effects. The adjusted Rand index values for the five scenarios are reported in Table 4. For all scenarios, the index is close to $1$ and confirms that our proposed model can correctly identify the principal strata and, combined with the good missing data imputation, allows us to estimate the expected associative and dissociative effects. Table 4 shows that our model correctly identifies the number of principal strata—two in Scenario 1 and 3, three in the others—while also providing accurate causal effect estimates. In contrast, neither the model by Lu et al. (2023) nor that of Schwartz et al. (2011)identify the principal strata according to our definition.
[…]
@@
Adjusted rand index for the five simulated scenarios computed on the point estimated partitions obtained with our proposed model. Scenario 1 Scenario 2 Scenario 3 Scenario 4 Scenario 5 Mean 0.9850 0.9906 0.9706 0.9717 0.9154 Standard deviation 0.0997 0.0597 0.1021 0.0892 0.1577
[…]
True and estimated principal causal effects—mean and standard deviation in the brackets—from our proposed method across scenarios. Scenario 1 Scenario 2 Scenario 3 Scenario 4 Scenario 5 True Estimated True Estimated True Estimated True Estimated True Estimated $\tau_{-}$ NA $-1.00$ $-0.98$ (0.32) NA $1.50$ $1.46$ (0.33) $-1.00$ $-0.79$ (0.68) $\tau_{0}$ $0.50$ $0.11$ (0.56) $4.00$ $4.03$ (0.61) $5.60$ $4.98$ (0.52) $4.00$ $4.11$ (0.53) $4.00$ $4.57$ (1.50) $\tau_{+}$ $9.00$ $8.79$ (1.25) $9.00$ $9.02$ (0.45) $8.60$ $9.36$ (0.87) $6.50$ $6.59$ (0.35) $9.00$ $8.91$ (0.44) {tabnote} NA: Not applicable. The stratum was not present in the scenario.